Given a die, it has 6 faces in which each face has either dot/s of π‘₯=1,2,3,4,5 π‘Žπ‘›π‘‘ 6. Given it as the population, consider sample of size 𝑛=3. Find the population mean, population variance, population standard deviation, the mean, variance and standard deviation of the sampling distribution of the sample mean and illustrate its probability histogram of the sampling distribution of the sample means.


1
Expert's answer
2022-04-16T04:12:57-0400

"\mu =\frac{1+2+3+4+5+6}{6}=3.5\\\sigma ^2=\frac{\left( 1-3.5 \right) ^2+\left( 2-3.5 \right) ^2+\left( 3-3.5 \right) ^2+\left( 4-3.5 \right) ^2+\left( 5-3.5 \right) ^2+\left( 6-3.5 \right) ^2}{6}=\frac{35}{12}\\\sigma =\sqrt{\sigma ^2}=\sqrt{\frac{35}{12}}\\\mu _{\bar{x}}=\mu =3.5\\{\sigma _{\bar{x}}}^2=\frac{N-n}{n\left( N-1 \right)}\sigma ^2=\frac{6-3}{3\left( 6-1 \right)}\cdot \frac{35}{12}=\frac{7}{12}\\\sigma _{\bar{x}}=\sqrt{{\sigma _{\bar{x}}}^2}=\sqrt{\frac{7}{12}}"